Full moon Corporation expects an EBIT of $26,000 every year forever. The company currently has no debt, and its cost of equity is 15 percent. The corporate tax rate is 35 percent.

What is the current value of the company?

Suppose the company can barrow at 10 percent. What will the value of the firm be if the company takes on debt equal to 40 percent of its unlevered value?

Suppose the company can barrow at 10 percent. What will the value of the firm be if the company takes on debt equal to 100 percent of its unlevered value?

What will the value of the firm be if the company takes on debt equal to 40 percent of its levered value?

What will the value of the firm be if the company takes on debt equal to 100 percent of its levered value?
Answer #1

Answer a.

Value of Unlevered Firm, VU = EBIT * (1 - tax) / Cost of Equity
Value of Unlevered Firm, VU = $26,000 * (1 - 0.35) / 0.15
Value of Unlevered Firm, VU = $112,666.67

Answer b.

Value of Debt, D = 40% * VU
Value of Debt, D = 40% * $112,666.67
Value of Debt, D = $45,066.67

Value of Levered Firm, VL = VU + tax * D
Value of Levered Firm, VL = $112,666.67 + 0.35 * $45,066.67
Value of Levered Firm, VL = $128,440.00

Answer c.

Value of Debt, D = 100% * VU
Value of Debt, D = 100% * $112,666.67
Value of Debt, D = $112,666.67

Value of Levered Firm, VL = VU + tax * D
Value of Levered Firm, VL = $112,666.67 + 0.35 * $112,666.67
Value of Levered Firm, VL = $152,100.00

Answer d.

Value of Debt, D = 40% * VL

Value of Levered Firm, VL = VU + tax * D
VL = $112,666.67 + 0.35 * 40% * VL
VL = $112,666.67 + 0.14 * VL
0.86 * VL = $112,666.67
VL = $131,007.76

Answer e.

Value of Debt, D = 100% * VL
Value of Debt, D = VL

Value of Levered Firm, VL = VU + tax * D
VL = $112,666.67 + 0.35 * VL
0.65 * VL = $112,666.67
VL = $173,333.34